OpenStudy (mrs.ambrose614):

The functions p(x) and g(x) are shown below: g(x) = 0.09x p(x) = (0.09)x Which statement best describes the graph of p(x) and g(x)? @pooja195 @satellite73 @zepdrix can you help me???

satellite73 (satellite73):

they are identical the way you have them written is one of them perhaps \((.09)^x\)?

OpenStudy (mrs.ambrose614):

Are you still their??

satellite73 (satellite73):

oui

OpenStudy (mrs.ambrose614):

Question 1 (Multiple Choice Worth 1 points) (04.05) The population f(x), in millions, of State A of a country after x years is represented by the function shown below: f(x) = 4(1.08)x The graph shows the population g(x), in millions, of State B of the country after x years: Graph of function g of x equals 2 multiplied by 1.08 to the power of x Which conclusion is correct about the population of State A and State B? The original population of State A was double of the original population of State B. The original population of State B was double of the original population of State A. The original population of State A was four times of the original population of State B. The original population of State A was equal to the original population of State B.

satellite73 (satellite73):

\[f(x) = 4(1.08)^x \\ g(x)=2(1..08)^x\] like that?

OpenStudy (mrs.ambrose614):

Yes

satellite73 (satellite73):

the original population of A is the 4 out front the original population of B is the 2 out front

OpenStudy (mrs.ambrose614):

So what's the awnser???

satellite73 (satellite73):

lol is 4 double 2?

OpenStudy (mrs.ambrose614):

What's!!!????????

satellite73 (satellite73):

here are your choices The original population of State A was double of the original population of State B. The original population of State B was double of the original population of State A. The original population of State A was four times of the original population of State B. The original population of State A was equal to the original population of State B.

satellite73 (satellite73):

4 was double 2 2 was double 4. 4was four times 2 4was equal to 2

satellite73 (satellite73):

which one of those statements is true?

OpenStudy (mrs.ambrose614):

I'd it b???

satellite73 (satellite73):

2 is double 4???

satellite73 (satellite73):

when i double 4, i get 8, not 2

OpenStudy (mrs.ambrose614):

So it would be c???

satellite73 (satellite73):

sure why not, if \[4=4\times 2\]

OpenStudy (mrs.ambrose614):

So is c correct???

OpenStudy (mathmale):

I suggest it would do you more good in the long run to defend your answer (explain your reasoning) than to depend on someone else to tell you whether you're right nor not.

OpenStudy (mrs.ambrose614):

He was explaining and stop giving negative comments he's allowed to help me if I am confused

OpenStudy (mathmale):

I think you've missed my point completely. Any independence you can develop will help you in the long run. You won't have satellite73 or anyone else with you when you take your next test. It'd be so great if you knew how to judge whether your own answers make sense. There is NOTHING negative about what I said to you last time.